5. Selected operating data for two divisions of Ashley Company is given below:

NORTH SOUTH

Sales 720,000 600,000

Total Assets 360,000 200,000

Fixed Assets 140,000 150,000

Total Liabilities 120,000 140,000

Long-Term Liabilities 80,000 20,000

Net income 72,000 48,000

Calculate ROA, ROE, Profit Margin and Fixed Assets Turnover for both the North and South divisions.
Answer #1

ROA = Net Income / Total Assets

North:

= 72,000 / 360,000

= 0.20 i.e 20%

South:

= 48,000 / 200,000

= 0.20 i.e 20%

ROE = Net Income / [ Total Assets - Total Liabilities ]

North :

= 72,000 / [ 360,000 - 120,000 ]

= 72,000 / 240,000

= 0.3 i.e 30%

South:

= 48,000 / [ 200,000 - 140,000 ]

= 48,000 / 60,000

= 0.8 i.e 80%

Profit Margin = Net Income / Sales

North:

= 72,000 / 720,000

= 0.10 i.e 10%

South:

48,000 / 600,000

= 0.08 i.e 8%

Fixed Asset Turnover = Sales / Fixed Assets

North:

= 720,000 / 140,000

= 5.14

South:

= 600,000 / 150,000

= 4

Pls comment, if any further assistance is required